# Stability of a passive viscous droplet in a confined active nematic liquid crystal

### v1

@@ III.2 Domain perturbation and linearized nematic field
Typical director fields corresponding to $m=1$ and $5$ are illustrated in Fig. 1(b,c).
@@ III.3 Linearized active and elastic stresses
$\bm{\sigma}^{e}=\bm{\sigma}^{e}_{0}+\epsilon\bm{\sigma}^{e}_{1}+O(\epsilon^{2})=-\frac{\xi^{2}}{r^{2}}\hat{\mathbf{r}}\hat{\mathbf{r}}-\epsilon\frac{\xi^{2}}{r^{2}}\left[2\partial_{\theta}\beta_{1}\hat{\mathbf{r}}\hat{\mathbf{r}}+\beta_{1}\big{(}\hat{\mathbf{r}}\hat{\bm{\theta}}+\hat{\bm{\theta}}\hat{\mathbf{r}}\big{)}\right]+O(\epsilon^{2}).$ (31)
@@ III.4 Linearized flow problem and boundary conditions
Our method of solution, described in Sec. III.5, will rely on the Lorentz reciprocal theorem for Stokes flow Masoud and Stone (2019). To this end, it is convenient to first recast the flow problem around a circular boundary ($r=\eta$). Note that Eqs. (32)–(33) remain unchanged around the circle to linear order. We also need to apply the dynamic boundary condition (17) on the circle, which is achieved by expanding the various fields involved in the boundary conditions as Taylor series in the neighborhood of $\zeta=0$. For example, the stress tensor can be expanded as:
@@
Inserting Eq. (34) into Eq. (17), linearizing, and simplifying using Eqs. (19)–(22), we obtain the following dynamic boundary condition at $O(\epsilon)$,
$\hat{\mathbf{r}}\cdot\big{(}\bm{\sigma}^{h}_{1}+\bm{\sigma}^{a}_{1}+\bm{\sigma}^{e}_{1}-\overline{\bm{\sigma}}^{h}_{1}\big{)}=\frac{\kappa_{1}\hat{\mathbf{r}}}{\mathrm{Ca}}+\frac{\widetilde{\zeta_{1}}}{\eta}\left(S-\frac{\xi^{2}}{\eta^{2}}\right)\big{(}\cos m\theta\,\hat{\mathbf{r}}+m\sin m\theta\,\hat{\bm{\theta}}\big{)}\qquad\mathrm{at}\,\,\,r=\eta.$ (35)
@@ III.5 Lorentz reciprocal theorem
where $\Sigma_{0}$ is an arbitrary constant. In two dimensions, this Newtonian problem is readily solved in terms of two streamfunctions $\psi(r,\theta)$ and $\overline{\psi}(r,\theta)$ such that
@@
with similar expressions for the streamfunction and its derivatives inside the drop. The two dimensionless functions $G_{m}(r)$ and $\overline{G}_{m}(r)$ can be obtained analytically by the solution of the Stokes equations; details of their solutions are presented in the Appendix.
@@
$\int_{\mathcal{C}_{0}}\hat{\mathbf{r}}\cdot\big{(}\bm{\sigma}^{h}_{1}+\bm{\sigma}^{e}_{1}+\bm{\sigma}^{a}_{1}-\overline{\bm{\sigma}}^{h}_{1}\big{)}\cdot\mathbf{U}\,\mathrm{d}\ell$ $=\int_{0}^{2\pi}\bigg{[}\frac{\kappa_{1}U_{r}}{\mathrm{Ca}}+\frac{\widetilde{\zeta}_{1}}{\eta}\left(S-\frac{\xi^{2}}{\eta^{2}}\right)\big{(}U_{r}\cos m\theta+U_{\theta}m\sin m\theta\big{)}\bigg{]}\eta\,\mathrm{d}\theta$ (47)
$=m\pi\Sigma_{0}\widetilde{\zeta}_{1}\left[\frac{m^{2}-1}{\mathrm{Ca}}\frac{G_{m}(\eta)}{\eta^{2}}+\left(S-\frac{\xi^{2}}{\eta^{2}}\right)\left(\frac{G_{m}(\eta)}{\eta}-G_{m}^{\prime}(\eta)\right)\right].$ (48)
@@
$\int_{\Omega_{0}^{+}}\big{(}\bm{\sigma}^{e}_{1}+\bm{\sigma}^{a}_{1}\big{)}:\mathbf{E}\,\mathrm{d}S$ $=\int_{0}^{2\pi}\int_{\eta}^{1}\bigg{[}2S\beta_{1}E_{r\theta}-\frac{2\xi^{2}}{r^{2}}\big{(}\partial_{\theta}\beta_{1}E_{rr}+\beta_{1}E_{r\theta}\big{)}\bigg{]}r\mathrm{d}r\mathrm{d}\theta$ (49)
$\begin{split}&=-\pi\Sigma_{0}\widetilde{\zeta}_{1}\frac{m\eta^{m-1}}{\eta^{2m}-1}\left\{S\int_{\eta}^{1}\big{(}r^{m}-r^{-m}\big{)}\bigg{[}rG_{m}^{\prime\prime}(r)-G_{m}^{\prime}(r)+m^{2}\frac{G_{m}(r)}{r}\bigg{]}\mathrm{d}r\right.\\ &\,\,\,\,\,\,-\xi^{2}\left.\int_{\eta}^{1}\big{(}r^{m}-r^{-m}\big{)}\bigg{[}\frac{G_{m}^{\prime\prime}(r)}{r}-(2m^{2}+1)\frac{G_{m}^{\prime}(r)}{r^{2}}+3m^{2}\frac{G_{m}(r)}{r^{3}}\bigg{]}\mathrm{d}r\right\}.\end{split}$ (50)
@@ III.6 Dispersion relation and growth rates
$\Xi^{a}_{m}(\eta,\lambda)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-m\frac{\mathrm{d}}{\mathrm{d}\eta}\bigg{[}\frac{G_{m}(\eta)}{\eta}\bigg{]}}-\frac{m\eta^{m-2}}{\eta^{2m}-1}\int_{\eta}^{1}\big{(}r^{m}-r^{-m}\big{)}\bigg{[}rG_{m}^{\prime\prime}(r)-G_{m}^{\prime}(r)+m^{2}\frac{G_{m}(r)}{r}\bigg{]}\mathrm{d}r,$ (54)
$\Xi^{e}_{m}(\eta,\lambda)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{m}{\eta^{2}}\frac{\mathrm{d}}{\mathrm{d}\eta}\bigg{[}\frac{G_{m}(\eta)}{\eta}\bigg{]}}+\frac{m\eta^{m-2}}{\eta^{2m}-1}\int_{\eta}^{1}\big{(}r^{m}-r^{-m}\big{)}\bigg{[}\frac{G_{m}^{\prime\prime}(r)}{r}-(2m^{2}+1)\frac{G_{m}^{\prime}(r)}{r^{2}}+3m^{2}\frac{G_{m}(r)}{r^{3}}\bigg{]}\mathrm{d}r.$ (55)
@@ III.6.1 Translation mode $(m=1)$
$\Xi^{e}_{1}(\eta,\lambda)$ $=\frac{2}{\eta}\,\overline{A}_{1}-\left[\frac{1}{\eta^{3}}+\frac{2\ln\eta}{\eta(1-\eta^{2})}\right]B_{1}+\frac{1-\eta^{2}}{\eta^{5}}D_{1},$ (57)
@@
Figure 2: (a) Active contribution $\Xi^{a}_{1}$ and (b) elastic contribution $\Xi^{e}_{1}$ to the growth rate $\Lambda_{1}$ for the translational mode ($m=1$), plotted as functions of $\eta$ for different viscosity ratios $\lambda$. (c) Color map of the total growth rate $\Lambda_{1}$ in the $(\eta,\lambda)$ parameter space for an extensile system ($S=+1$) with $\xi^{2}=1$. Linearly stable and unstable regions are shown in blue and red, respectively, and the grey curve shows the marginal stability curve. (d) Effect of $\xi^{2}$ on the marginal stability curve in (c). Panels (e) and (f) show the same as (c) and (d) for a contractile system ($S=-1$).
@@
The elastic contribution shows a more complex dependence on both $\eta$ and $\lambda$ in Fig. 2(b). High-viscosity drops are always stable under elastic stresses. As $\lambda$ decreases, an unstable region appears at intermediate values of $\eta$. Regardless of $\lambda$, we find that $\Xi^{e}_{1}\to-\infty$ as $\eta\to 0$, i.e., very small drops are always stable to translation. The divergence of the growth rate in that limit is attributed to the divergence of the base-state elastic pressure at the origin when $\eta\to 0$ as seen in Eq. (21). When both active and elastic stresses play a role, the translational mode can be either stable or unstable depending on parameter values. For the extensile case, this is illustrated in Fig. 2(c), showing a color plot of the total growth rate $\Lambda_{1}$ in the $(\eta,\lambda)$ parameter space: the system is primarily stable, except in the red region which occurs at intermediate values of $\eta$ and low viscosity ratios. The dependence of the marginal stability curve on $\xi^{2}$ is shown in Fig. 2(d), where increasing the magnitude of elastic stresses is found to increase the size of the unstable region. The contractile case is analyzed similarly in Fig. 2(e,f): the unstable region is found to be larger in that case, but to decrease in size with increasing $\xi^{2}$.
@@
$f^{e}_{r1}=-(\eta^{2}-1)^{-1}(r^{-2}-3r^{-4})\cos\theta,\qquad f^{e}_{\theta 1}=(\eta^{2}-1)^{-1}(r^{-2}-r^{-4})\sin\theta.$ (60)
The two force fields are shown in Fig. 3 for a representative case with $\eta=0.45$, for a circular drop that was displaced to the right from its equilibrium position; note that the active force field is for an extensile system and changes sign in the contractile case. The active and elastic force bulk densities display a complex structure and are strongest near the drop surface, where they would appear to have a destabilizing effect by pushing the fluid and drop further to the right. It should be kept in mind, however, that these force fields are applied on the solvent in the nematic layer, where they drive a flow whose structure is not explicitly known as the Lorentz reciprocal theorem circumvented its calculation. First-order active and elastic tractions on the interface, i.e., $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{a,e}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{a,e}$, are shown in the lower-right panels in Fig. 3, where they both appear to resist translation of the drop and thus have a stabilizing effect. Here again, these traction fields do not account for the additional hydrodynamic tractions resulting from flow fields inside and outside the drop, and thus provide an incomplete description of stresses acting on the interface. Nevertheless, the stresses in Fig. 3 suggest a subtle interplay of active, elastic and viscous effects in governing the growth rates calculated in Fig. 2.
Figure 3: (a) Active and (b) elastic bulk force densities and surface tractions at linear order in $\epsilon$ for the translation mode $m=1$, in the case of a drop with radius $\eta=0.45$. In each case, the left plot shows the first-order force density, $\mathbf{f}^{a,e}_{1}=\nabla\cdot\bm{\sigma}^{a,e}_{1}$, in the bulk of the nematic, while the right plot shows the first-order traction field, $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{a,e}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{a,e}$, on the drop surface. Panel (a) shows the active force and traction for an extensile system with $S=+1$.

### v2

@@ III.2 Domain perturbation and linearized nematic field
Typical director fields corresponding to $m=1$ and 555 are illustrated in Fig. 1(b,c).
@@ III.3 Linearized active and elastic stresses
$\bm{\sigma}^{e}=\bm{\sigma}^{e}_{0}+\epsilon\bm{\sigma}^{e}_{1}+O(\epsilon^{2})=-\frac{\xi^{2}}{r^{2}}\hat{\bm{\theta}}\hat{\bm{\theta}}-\epsilon\frac{\xi^{2}}{r^{2}}\left[2\partial_{\theta}\beta_{1}\hat{\bm{\theta}}\hat{\bm{\theta}}+r\partial_{r}\beta_{1}\big{(}\hat{\mathbf{r}}\hat{\bm{\theta}}+\hat{\bm{\theta}}\hat{\mathbf{r}}\big{)}\right]+O(\epsilon^{2}).$ (32)
@@ III.4 Linearized flow problem and boundary conditions
Our method of solution, described in Sec. III.5, will rely on the Lorentz reciprocal theorem for Stokes flow Masoud and Stone (2019). To this end, it is convenient to first recast the flow problem around a circular boundary ($r=\eta$). Specifically, we need to apply the dynamic boundary condition (18) on the circle, which is achieved by expanding the relevant fields as Taylor series in the neighborhood of $\zeta=0$. For example, the stress tensors can be expanded as:
@@
The elastic surface traction, which arises at order $\epsilon$, is expanded separately as
$\begin{split}\mathbf{t}^{e}_{s,1}=\epsilon\frac{\xi^{2}}{\eta}\bigg{[}\left(\partial_{\theta}\partial_{r}\beta_{1}|_{r=\eta}+\widetilde{\zeta_{1}}\frac{m^{2}}{\eta^{2}}\cos m\theta\right)\hat{\mathbf{r}}+\left(\partial_{r}\beta_{1}|_{r=\eta}+\widetilde{\zeta_{1}}\frac{m}{\eta^{2}}\sin m\theta\right)\hat{\bm{\theta}}\bigg{]}+O(\epsilon^{2}).\end{split}$ (36)
Inserting Eqs. (35) and (36) into Eq. (18), linearizing, and simplifying using Eqs. (20)–(23), we obtain the following dynamic boundary condition at $O(\epsilon)$,
$\begin{split}\hat{\mathbf{r}}\cdot\big{(}\bm{\sigma}^{h}_{1}+\bm{\sigma}^{a}_{1}+\bm{\sigma}^{e}_{1}-\overline{\bm{\sigma}}^{h}_{1}\big{)}&=\frac{\kappa_{1}\hat{\mathbf{r}}}{\mathrm{Ca}}+\frac{\widetilde{\zeta_{1}}}{\eta}S\big{(}\cos m\theta\,\hat{\mathbf{r}}+m\sin m\theta\,\hat{\bm{\theta}}\big{)}\\ &-\widetilde{\zeta_{1}}\frac{\xi^{2}}{\eta^{3}}\bigg{[}\bigg{(}\frac{m^{3}(\eta^{2m}+1)}{\eta^{2m}-1}+m^{2}-1\bigg{)}\cos m\theta\,\hat{\mathbf{r}}+\frac{m^{2}(\eta^{2m}+1)}{\eta^{2m}-1}\sin m\theta\,\hat{\bm{\theta}}\bigg{]}\qquad\mathrm{at}\,\,\,r=\eta.\end{split}$ (37)
@@ III.5 Lorentz reciprocal theorem
where $\Sigma_{0}$ is an arbitrary constant. In two dimensions, this Newtonian fluid problem is readily solved in terms of two streamfunctions $\psi(r,\theta)$ and $\overline{\psi}(r,\theta)$ such that
@@
with similar expressions for the streamfunction and its derivatives inside the drop. The two dimensionless functions $G_{m}(r)$ and $\overline{G}_{m}(r)$ are obtained analytically by solving the Stokes equations; details are presented in the Appendix.
@@
$\int_{\mathcal{C}_{0}}\hat{\mathbf{r}}\cdot\big{(}\bm{\sigma}^{h}_{1}+\bm{\sigma}^{e}_{1}+\bm{\sigma}^{a}_{1}-\overline{\bm{\sigma}}^{h}_{1}\big{)}\cdot\mathbf{U}\,\mathrm{d}\ell$ $=\int_{0}^{2\pi}\bigg{\{}\frac{\kappa_{1}U_{r}}{\mathrm{Ca}}+\frac{\widetilde{\zeta}_{1}S}{\eta}\big{(}U_{r}\cos m\theta+U_{\theta}m\sin m\theta\big{)}$ (49)
$\,\,\,\,\,\,-\widetilde{\zeta_{1}}\frac{\xi^{2}}{\eta^{3}}\bigg{[}\bigg{(}\frac{m^{3}(\eta^{2m}+1)}{\eta^{2m}-1}+m^{2}-1\bigg{)}U_{r}\cos m\theta+\frac{m^{2}(\eta^{2m}+1)}{\eta^{2m}-1}U_{\theta}\sin m\theta\bigg{]}\bigg{\}}\eta\,\mathrm{d}\theta$
$\begin{split}&=m\pi\Sigma_{0}\widetilde{\zeta}_{1}\bigg{\{}\left[\frac{m^{2}-1}{\mathrm{Ca}}\frac{G_{m}(\eta)}{\eta^{2}}-S\left(G_{m}^{\prime}(\eta)-\frac{G_{m}(\eta)}{\eta}\right)\right]\\ &\,\,\,\,\,\,+\frac{\xi^{2}}{\eta^{2}}\bigg{[}\frac{m(\eta^{2m}+1)}{\eta^{2m}-1}G^{\prime}_{m}(\eta)-\bigg{(}\frac{m^{3}(\eta^{2m}+1)}{\eta^{2m}-1}+m^{2}-1\bigg{)}\frac{G_{m}(\eta)}{\eta}\bigg{]}\bigg{\}}.\end{split}$ (50)
@@
$\int_{\Omega_{0}^{+}}\big{(}\bm{\sigma}^{e}_{1}+\bm{\sigma}^{a}_{1}\big{)}:\mathbf{E}\,\mathrm{d}S$ $=\int_{0}^{2\pi}\int_{\eta}^{1}\bigg{[}2S\beta_{1}E_{r\theta}-\frac{2\xi^{2}}{r^{2}}\big{(}\partial_{\theta}\beta_{1}E_{\theta\theta}+r\partial_{r}\beta_{1}E_{r\theta}\big{)}\bigg{]}r\mathrm{d}r\mathrm{d}\theta$ (51)
$\begin{split}&=-\pi\Sigma_{0}\widetilde{\zeta}_{1}\frac{m\eta^{m-1}}{\eta^{2m}-1}\left\{S\int_{\eta}^{1}\big{(}r^{m}-r^{-m}\big{)}\bigg{[}rG_{m}^{\prime\prime}(r)-G_{m}^{\prime}(r)+m^{2}\frac{G_{m}(r)}{r}\bigg{]}\mathrm{d}r\right.\\ &\,\,\,\,\,\,-2m^{2}\xi^{2}\int_{\eta}^{1}\big{(}r^{m-2}-r^{-m-2}\big{)}\bigg{[}G_{m}^{\prime}(r)-\frac{G_{m}(r)}{r}\bigg{]}\mathrm{d}r\\ &\,\,\,\,\,\,-m\xi^{2}\left.\int_{\eta}^{1}\big{(}r^{m-2}+r^{-m-2}\big{)}\bigg{[}rG_{m}^{\prime\prime}(r)-G_{m}^{\prime}(r)+m^{2}\frac{G_{m}(r)}{r}\bigg{]}\mathrm{d}r\right\}.\end{split}$ (52)
@@ III.6 Dispersion relation and growth rates
$\Xi^{a}_{m}(\eta,\lambda)$ $=-m\frac{\mathrm{d}}{\mathrm{d}\eta}\bigg{[}\frac{G_{m}(\eta)}{\eta}\bigg{]}-\frac{m\eta^{m-2}}{\eta^{2m}-1}\int_{\eta}^{1}\big{(}r^{m}-r^{-m}\big{)}\bigg{[}rG_{m}^{\prime\prime}(r)-G_{m}^{\prime}(r)+m^{2}\frac{G_{m}(r)}{r}\bigg{]}\mathrm{d}r,$ (56)
$\begin{split}\Xi^{e}_{m}(\eta,\lambda)&={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{m}{\eta^{3}}\bigg{[}\frac{m(\eta^{2m}+1)}{\eta^{2m}-1}G^{\prime}_{m}(\eta)-\bigg{(}\frac{m^{3}(\eta^{2m}+1)}{\eta^{2m}-1}+m^{2}-1\bigg{)}\frac{G_{m}(\eta)}{\eta}\bigg{]}}\\ &\,\,\,\,\,\,\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\frac{m\eta^{m-2}}{\eta^{2m}-1}\bigg{\{}2m^{2}\int_{\eta}^{1}\big{(}r^{m-2}-r^{-m-2}\big{)}\bigg{[}G_{m}^{\prime}(r)-\frac{G_{m}(r)}{r}\bigg{]}\mathrm{d}r}\\ &\,\,\,\,\,\,\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+m\int_{\eta}^{1}\big{(}r^{m-2}+r^{-m-2}\big{)}\bigg{[}rG_{m}^{\prime\prime}(r)-G_{m}^{\prime}(r)+m^{2}\frac{G_{m}(r)}{r}\bigg{]}\mathrm{d}r\bigg{\}}}.\end{split}$ (57)
@@ III.6.1 Translation mode $(m=1)$
$\Xi^{e}_{1}(\eta,\lambda)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-2\bigg{(}\frac{\eta^{2}+1}{\eta^{2}-1}\bigg{)}\,\frac{\overline{A}_{1}}{\eta}-\frac{4}{\eta}A_{1}+\left[\frac{1}{\eta^{3}}+\frac{2\ln\eta}{\eta(1-\eta^{2})}\right]B_{1}-2\frac{1+\eta^{2}}{\eta^{5}}D_{1}},$ (59)
@@
Figure 2: (a) Active contribution $\Xi^{a}_{1}$ and (b) elastic contribution $\Xi^{e}_{1}$ to the growth rate $\Lambda_{1}$ for the translational mode ($m=1$), plotted as functions of $\eta$ for different viscosity ratios $\lambda$. (c) Color map of the total growth rate $\Lambda_{1}$ in the $(\eta,\lambda)$ parameter space for an extensile system ($S=+1$) with $\xi^{2}=1$. Linearly stable (S) and unstable (U) regions are shown in blue and red, respectively, and the green curve shows the marginal stability curve. (d) Effect of $\xi^{2}$ on the marginal stability curve in (c). Panels (e) and (f) show the same as (c) and (d) for a contractile system ($S=-1$).
@@
The elastic contribution shows a more complex dependence on both $\eta$ and $\lambda$ in Fig. 2(b). Larger drops are always stable under elastic stresses, especially at high viscosity ratios. Regardless of $\lambda$, we find, however, that $\Xi^{e}_{1}\to+\infty$ as $\eta\to 0$, i.e., small drops are always unstable to translation. The divergence of the growth rate in that limit is attributed to the divergence of the base-state elastic pressure at the origin when $\eta\to 0$ as seen in Eq. (22). When both active and elastic stresses play a role, the translational mode can be either stable or unstable depending on parameter values. For the extensile case, this is illustrated in Fig. 2(c), showing a color plot of the total growth rate $\Lambda_{1}$ in the $(\eta,\lambda)$ parameter space: the system is always unstable for small drops and becomes stable as $\eta$ is increased, with the extent of the unstable region shrinking with increasing viscosity ratio. The dependence of the marginal stability curve on $\xi^{2}$ is shown in Fig. 2(d), where increasing the magnitude of elastic stresses is found to increase the size of the unstable region. The contractile case is analyzed similarly in Fig. 2(e,f): the unstable region is found to be larger in that case, but to decrease in size with increasing $\xi^{2}$.
@@
$f^{e}_{r1}=(\eta^{2}-1)^{-1}(r^{-2}-3r^{-4})\cos\theta,\qquad f^{e}_{\theta 1}=(\eta^{2}-1)^{-1}(r^{-2}-r^{-4})\sin\theta.$ (62)
The two force fields are shown in Fig. 3 for a representative case with $\eta=0.45$, for a circular drop that was displaced to the right from its equilibrium position; the active force field shown here is for an extensile system and changes sign in the contractile case. The active and elastic force bulk densities display a complex structure and are strongest near the drop surface, where they would appear to have a destabilizing effect by pushing the fluid and drop further to the right. It should be kept in mind, however, that these force fields are applied on the solvent in the nematic layer, where they drive a flow whose structure is not explicitly known as the Lorentz reciprocal theorem circumvented its calculation. First-order active and elastic tractions on the interface, $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{a}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{a}$ and $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{e}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{e}+\mathbf{t}_{s,1}^{e}$, respectively, are shown in the lower-right panels in Fig. 3, where they both appear to resist translation of the drop and thus have a stabilizing effect. Here again, these traction fields do not account for the additional hydrodynamic tractions resulting from flow fields inside and outside the drop, and thus provide an incomplete description of stresses acting on the interface. Nevertheless, the stresses in Fig. 3 suggest a subtle interplay of active, elastic and viscous effects in governing the growth rates calculated in Fig. 2.
Figure 3: (a) Active and (b) elastic bulk force densities and surface tractions at linear order in $\epsilon$ for the translation mode $m=1$, in the case of a drop with radius $\eta=0.45$. In each case, the left plot shows the first-order force density, $\mathbf{f}^{a,e}_{1}=\nabla\cdot\bm{\sigma}^{a,e}_{1}$, in the bulk of the nematic, while the right plot shows the first-order traction field, $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{a}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{a}$ or $\mathbf{n}_{0}\cdot\bm{\sigma}_{1}^{e}+\mathbf{n}_{1}\cdot\bm{\sigma}_{0}^{e}+\mathbf{t}_{s,1}^{e}$, on the drop surface. Panel (a) shows the active force and traction for an extensile system with $S=+1$.